# Compositional Risk Minimization

### v1

@@ 4.3 Extrapolation of Discriminative Model
A complete proof is provided in the Appendix. Observe that $\hat{p}(\cdot|x)=p(\cdot|x)$ is a condition that even a model trained via ERM can satisfy (with sufficient capacity and data) but it cannot match the true $q(\cdot|x)$. In contrast, CRM optimally adjusts the additive-energy classifier for the compositional shifts. CRM requires the knowledge of $q(\cdot)$ but the choice of uniform distribution over all possible groups is a safe one to make in the absence of knowledge of $q(\cdot)$. Notice how learned bias $\hat{B}(z)$ can only be fitted for $z\in\mathcal{Z}^{\mathrm{train}}$, remaining undefined for $z^{\prime}\notin\mathcal{Z}^{\mathrm{train}}$. But we can compute the extrapolated bias $B^{\star}(z^{\prime})$, $\forall z^{\prime}\in\mathcal{Z}^{\mathrm{test}}$, based remarkably on only data from the training distribution.
In the discussion so far, we have relied on a crucial assumption that the attribute combinations in the test distribution are in the affine hull. Is this also a necessary condition? Can we generalize to attributes outside the affine hull? We consider the task of learning $p(\cdot|z)$ from Theorem 1 and the task of learning $q(\cdot|x)$ from Theorem 2. In Section 9.5 in the Appendix, we show that the restriction to affine hulls is indeed necessary.
@@
Under the assumption of compositional shifts, we know that the support of $q(z)$, $\mathcal{Z}^{\mathsf{test}}$ is only restricted to be a subset of the Cartesian product set ${\mathcal{Z}^{\times}}$, but our results so far have required us to restrict the support further by confining it to the affine hull, i.e., $\mathcal{Z}^{\mathsf{test}}\subseteq\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})\subseteq{\mathcal{Z}^{\times}}$. This leads us to a natural question. If the training groups that form $\mathcal{Z}^{\mathsf{train}}$ are drawn at random, then how many groups do we need such that the affine hull captures ${\mathcal{Z}^{\times}}$, i.e., $\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})={\mathcal{Z}^{\times}}$, at which point CRM can achieve Cartesian Product Extrapolation. Another way to think about this is to say, how fast does the affine hull grow and capture the Cartesian product set ${\mathcal{Z}^{\times}}$?
@@
Consider the the setting with $m=2$ attribute dimensions, where each attribute takes $d$ possible values. In such a case, we have $d^{2}$ possible attribute combinations. Suppose we sample $s$ attribute vectors $z$ that comprise the support $\mathcal{Z}^{\mathsf{train}}$ uniformly at random (with replacement) from these $d^{2}$ possibilites. In the next theorem, we show that if the number of sampled attribute vectors exceeds $8cd\log(d)$, then the affine hull of $\mathcal{Z}^{\mathsf{train}}$ contains all the possible $d^{2}$ combinations with a high probability and as a result CRM achieves CPE.
@@
Consider the setting where $p(.|z)$ follows AED $\forall z\in{\mathcal{Z}^{\times}}$, $\mathcal{Z}^{\mathsf{train}}$ comprises of $s$ attribute vectors $z$ drawn uniformly at random from ${\mathcal{Z}^{\times}}$, and the test distribution $q$ satisfies compositional shift characterization. If $s\geq 8cd\log(d/2)$, where $d$ is sufficiently large, $\hat{p}(z|x)=p(z|x),\forall z\in\mathcal{Z}^{\mathsf{train}},\forall x\in\mathbb{R}^{n}$, $\hat{q}(z)=q(z),\forall z\in\mathcal{Z}^{\mathsf{test}}$, then the output of CRM (equation 9) matches the test distribution, i.e., $\hat{q}(z|x)=q(z|x)$, $\forall z\in\mathcal{Z}^{\mathsf{test}},\forall x\in\mathbb{R}^{n}$, with probability greater than $1-\frac{1}{c}$.
For the more general setting of $m$ attributes, we conjecture that a polynomial growth in $md$, i.e., $\mathcal{O}(\mathsf{poly}(md))$, groups suffice to generalize to distributions whose support span $d^{m}$ groups. To support this conjecture, we conduct numerical experiments (described in the Appendix), where we show that a random $z^{\prime}\in{\mathcal{Z}^{\times}}$ is in the affine span of a random set of $\mathcal{O}(md)$ training groups $z$ with a high probability. To summarize, these results point to a surprising fact that, we need to see data from a much smaller number of groups to achieve extrapolation to an exponentially large set.
5 Algorithm for Compositional Risk Minimization (CRM)
In a nutshell, CRM consists of: a) training a model of the form of equation 6 by maximum likelihood (equation 8) for trainset group prediction; b) compute extrapolated biases (equation 10); c) infer group probabilities on compositionally shifted test distribution using equation 9. For the case where we have 2 attributes $z=(y,a)$, Figure 1 illustrates a basic architecture using a deep network backbone $\phi(x;\theta)$ followed by a linear mapping (matrix $W$), and Algorithm 1 provides the associated pseudo-code. [^1] The figure’s architecture computes the logits $F_{L,B}(x)$ as implemented in the pseudocode. Alternatively to a single linear head whose output we split, we could use separate arbitrary (non-linear) heads to obtain the components for each attribute. Architecture and code can easily be generalized to handle more than 2 attributes.
Input: training set $\mathcal{D}^{\mathrm{train}}$ with examples $(x,y,a)$, where $y$ is the class to predict and $a$ is an attribute spuriously correlated with $y$
Output: classifier parameters $\theta$, $W$, $B^{\star}$.
  • Let $L,B\in\mathbb{R}^{d_{y}\times d_{a}}$ be the log prior and the bias terms.
  • Define logits: $F_{L,B}(x):=-((W\cdot\phi(x;\theta))_{1:d_{y}}+(W\cdot\phi(x;\theta))_{d_{y}+1:d_{y}+d_{a}}^{\top})+L-B$
  • Define log probabilities: $\log p(y,a|x;\theta,W,L,B):=(F_{L,B}(x)-\mathrm{logsumexp}(F_{L,B}(x)))_{y,a}$
@@ 5 Algorithm for Compositional Risk Minimization (CRM)
  • Estimate log prior $L^{\mathrm{train}}$ from $\mathcal{D}^{\mathrm{train}}$; $L^{\mathrm{train}}_{y,a}\leftarrow-\infty$ if $(y,a)$ absent from $\mathcal{D}^{\mathrm{train}}$.
  • Optimize $\theta$, $W$, and $B$ to maximize the log-likelihood over $\mathcal{D}^{\mathrm{train}}$:
$\theta,W,B\leftarrow\arg\max_{\theta,W,B}\sum_{(x,y,a)\in\mathcal{D}^{\mathrm{train}}}\log p(y,a|x;\theta,W,L^{\mathrm{train}},B)$
  • Extrapolate bias: $B^{\star}\leftarrow\log\left(\frac{1}{n}\sum_{x\in\mathcal{D}^{\mathrm{train}}}\exp(F_{0,0}(x)-\mathrm{logsumexp}(F_{L^{\mathrm{train}},B}(x)))\right)$
@@
    • Compute group probabilities, using $B^{\star}$, and $L^{\mathrm{unif}}=\log\frac{1}{d_{y}d_{a}}$ aiming for shift to uniform prior:
    $q(y,a|x)\leftarrow\exp(\log p(y,a|x;\theta,W,L^{\mathrm{unif}},B^{\star}))$
  • Marginalize over $a$ to get class probabilities: $q(y|x)\leftarrow\sum_{a}q(y,a|x)$
@@ 6.1 Setup
    We evaluate CRM on widely recognized benchmarks for subpopulation shifts (Yang et al., 2023b), that have 2 attributes $z=(y,a)$, where $y$ denotes the class label and $a$ denotes the spurious attribute ($y$ and $a$ are correlated). However, the standard split between train and test data mandated in these benchmarks does not actually evaluate robustness to compositional shifts, because both train and test datasets contain all the groups ($\mathcal{Z}^{\mathsf{train}}=\mathcal{Z}^{\mathsf{test}}={\mathcal{Z}^{\times}}$). Therefore, we repurpose these benchmarks for compositional shifts by discarding samples from one of the groups ($z$) in the train (and validation) dataset; but we don’t change the test dataset, i.e., $z\not\in\mathcal{Z}^{\mathsf{train}}$ but $z\in\mathcal{Z}^{\mathsf{test}}$. Let us denote the data splits from the standard benchmarks as $({\mathcal{D}}_{\mathsf{train}},{\mathcal{D}}_{\mathsf{val}},{\mathcal{D}}_{\mathsf{test}})$. Then we generate multiple variants of compositional shifts $\{({\mathcal{D}}_{\mathsf{train}}^{\neg z},{\mathcal{D}}_{\mathsf{val}}^{\neg z},{\mathcal{D}}_{\mathsf{test}})\;|\;z\in{\mathcal{Z}^{\times}}\}$, where ${\mathcal{D}}_{\mathsf{train}}^{\neg z}$ and ${\mathcal{D}}_{\mathsf{val}}^{\neg z}$ are generated by discarding samples from ${\mathcal{D}}_{\mathsf{train}}$ and ${\mathcal{D}}_{\mathsf{val}}$ that belong to the group $z$.
@@
    Following this procedure, we adapted Waterbirds (Wah et al., 2011), CelebA (Liu et al., 2015), MetaShift (Liang and Zou, 2022), MultiNLI (Williams et al., 2017), and CivilComments Borkan et al. (2019) for experiments. We also experiment with the NICO++ dataset (Zhang et al., 2023), where we already have $\mathcal{Z}^{\mathsf{train}}\subsetneq\mathcal{Z}^{\mathsf{test}}={\mathcal{Z}^{\times}}$ as some groups were not present in the train dataset. However, these groups are still present in the validation dataset ($\mathcal{Z}^{\mathsf{val}}={\mathcal{Z}^{\times}}$). Hence, the only transformation we apply to NICO++ is to drop samples from the validation dataset such that $\mathcal{Z}^{\mathsf{train}}=\mathcal{Z}^{\mathsf{val}}$.
@@
    For baselines, we train classifiers via Empirical Risk Minimization (ERM), GroupDRO (Sagawa et al., 2019), Logit Correction (LC) (Liu et al., 2022b), and supervised logit adjustment (sLA) (Tsirigotis et al., 2024). In all cases we employ a pretrained architecture as the representation network $\phi$, followed by a linear layer $W$ to get class predictions, and fine-tune them jointly (see Section 10.3 for details). For evaluation metrics, we report the average accuracy, group balanced accuracy, and worst-group accuracy on the test dataset. Due to imbalances in group distribution, a method can obtain good average accuracy despite having bad worst-group accuracy. Therefore, the worst-group accuracy is a more indicative metric of robustness to spurious correlations (more details in Section 10.2).
@@ 6.2 Results
    Table 1 shows the results of our experiment. For each dataset, we report the average accuracy over its various compositional shift scenarios $\{({\mathcal{D}}_{\mathsf{train}}^{\neg z},{\mathcal{D}}_{\mathsf{val}}^{\neg z},{\mathcal{D}}_{\mathsf{test}})\;|\;z\in{\mathcal{Z}^{\times}}\}$ (detailed results for all scenarios are in Section 11.1). In all cases, CRM either outperforms or is competitive with the baselines in terms of worst group accuracy (WGA). Further, for Waterbirds and MultiNLI, while the logit adjustment baselines appear competitive with CRM on average, if we look more closely at the worst case compositional shift scenario, we find that logit adjustment baselines fare much worse than CRM. For Waterbirds, LC obtains $69.0\%$ worst group accuracy while CRM obtains $73.0\%$ worst group accuracy for the worst case scenario of dropping the group $(0,1)$ (Table 5). Similarly, for the MultiNLI benchmark, sLA obtains $19.7\%$ worst group accuracy while CRM obtains $31.0\%$ worst group accuracy for the worst case scenario of dropping the group $(0,0)$ (Table 8).
@@
    We also report the worst group accuracy (other metrics in Table 11) for the original benchmark $({\mathcal{D}}_{\mathsf{train}},{\mathcal{D}}_{\mathsf{val}},{\mathcal{D}}_{\mathsf{train}})$ which was not transformed for compositional shifts, denoted WGA (No Groups Dropped). WGA (No Groups Dropped) can be interpreted as the “oracle” performance for that benchmark, and we can compare methods based on the performance drop in WGA due to discarding groups in compositional shifts. ERM and GroupDRO appear the most sensitive to compositional shifts, and the logit adjustment baselines also show a sharp drop for the CelebA benchmark; while CRM is more robust to compositional shifts.
    Importance of extrapolating the bias.
@@
    We conduct an ablation study for CRM where we test a variant that uses the learned bias $\hat{B}$ (e.q. 8) instead of the extrapolated bias $B^{\star}$ (e.q. 10). Results are presented in Table 2. They show a significant drop in worst-group accuracy if we use the learned bias instead of the extrapolated one. Hence, our theoretically grounded bias extrapolation step is crucial to generalize under compositional shifts. In Section 11.1 (Table 10) we conduct further ablation studies, showing the impact of different choices of the test log prior.
    7 Conclusion
    We provide a novel approach based on flexible additive energy models for compositionality in discriminative tasks. Our proposed CRM approach can provably extrapolate to novel attribute combinations within the discrete affine hull of the training support, where the affine hull grows quickly with the training groups to cover the Cartesian product extension of the training support. Our empirical results demonstrate that the additive energy assumption is sufficiently flexible to yield good classifiers for high-dimensional images, and that the proposed CRM estimator is able to extrapolate to novel combinations in $\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})$, without having to model high-dimensional $p(x|z)$ nor having to estimate their partition function. CRM is a simple and efficient algorithm that empirically proved consistently more robust to compositional shifts than approaches based on other logit-shifting schemes and GroupDRO.
@@ 7 Conclusion
    Table 1: Robustness under compositional shift. We compare the proposed Compositional Risk Minimization (CRM) method to baseline Expected Risk Minimization (ERM) classifier training with no group information, and to robust methods that leverage group labels: Group-DRO (G-DRO) (Sagawa et al., 2019), Logit Correction (LC) (Liu et al., 2022b) and Supervised Logit Adjustment (sLA) (Tsirigotis et al., 2024). We report test Average Accuracy and Worst Group Accuracy (WGA), averaged as a group is dropped from training and validation sets. Last column is WGA under the dataset’s standard subpopulation shift benchmark, i.e. with no group dropped. All methods have a harder time to generalize when groups are absent from training, but CRM appears consistently more robust (standard error based on 3 random seeds).

### v2

@@ 2.3 Extrapolation of Discriminative Model
A complete proof is provided in Section 10.2. Observe that $\hat{p}(\cdot|x)=p(\cdot|x)$ is a condition that even a model trained via ERM can satisfy (with sufficient capacity and data) but it cannot match the true $q(\cdot|x)$. In contrast, CRM optimally adjusts the additive-energy classifier for the compositional shifts. CRM requires the knowledge of test prior $q(z)$ but the choice of uniform distribution over all possible groups is a reasonable one to make in the absence of further knowledge. Notice how learned bias $\hat{B}(z)$ can only be fitted for $z\in\mathcal{Z}^{\mathrm{train}}$, remaining undefined for $z^{\prime}\notin\mathcal{Z}^{\mathrm{train}}$. But we can compute the extrapolated bias $B^{\star}(z^{\prime})$, $\forall z^{\prime}\in\mathcal{Z}^{\mathrm{test}}$, based remarkably on only data from the train distribution.
Illustrating CRM’s adaptation to test distribution.
To better convey how CRM can adapt to the Bayes optimal classifier of the test distribution, we provide an example. Consider a two-dimensional setting, where the distribution of $x\in\mathbb{R}^{2}$ conditioned on the attributes $z_{1}\in\{-1,1\}$ and $z_{2}\in\{-1,1\}$ is a Gaussian with mean $(z_{1},z_{2})$ and identity covariance. Suppose the training groups are drawn with equal probability and can take one of the following three possible values $(+1,+1),(-1,+1),(+1,-1)$. We do not observe data from the group $(-1,-1)$ during training, but at test time we draw samples from all the four groups with equal probability. First, we can show that the above distribution can be expressed as an additive energy distribution, as $E(x,(z_{1},z_{2}))=\frac{1}{4}\|x-(2z_{1},0)\|^{2}+\frac{1}{4}\|x-(0,2z_{2})\|^{2})$. For the task of classifying groups, the Bayes optimal classifier has a closed form solution, where each decision region is an intersection of two half-spaces. Figure 2 shows that CRM learns the Bayes optimal classifier on the training distribution, enables shifting from train prior to test prior to yield the Bayes optimal classifier for the test distribution, and correctly generalize to the unseen $(-1-1)$ group. For further details, including illustration of the failure of ERM-trained binary classifier on this problem, see Section 11. Beyond this toy example, we highlight that the additive energy form supports modeling nearly arbitrarily complex distributions.
Figure 2: Extrapolating to an unseen test group. The distribution to model corresponds to a mixture of 4 Gaussians. But the $(-1,-1)$ group (pink dashed) has zero prior probability in the training distribution, i.e., is absent from the training set. As a result, discriminative training (left) learns only three decision regions and would misclassify test points from $(-1,-1)$. CRM adjusts the prior at test time (right) to a uniform distribution over all attribute combinations, enabling it to recover four decision regions and correctly generalize to the unseen $(-1,-1)$ group. The additive energy form composes the attributes’ information to yield this unseen group’s location. Decision regions were obtained from finite-data simulations, leading to minor imperfections.
Analyzing growth of Discrete Affine Hull.
@@
In the discussion so far, we have relied on a crucial assumption that the attribute combinations in the test distribution are in the affine hull. Is this also a necessary condition? Can we generalize to attributes outside the affine hull? We consider the task of learning $p(\cdot|z)$ from Theorem 1 and the task of learning $q(\cdot|x)$ from Theorem 2. In  Section 10.6, we show that the restriction to affine hulls is indeed necessary. Under the assumption of compositional shifts $\mathcal{Z}^{\mathsf{test}}$ is only restricted to be a subset of the Cartesian product set ${\mathcal{Z}^{\times}}$, but our results so far have required us to restrict the support further by confining it to the affine hull, i.e., $\mathcal{Z}^{\mathsf{test}}\subseteq\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})\subseteq{\mathcal{Z}^{\times}}$. This leads us to a natural question. If the training groups that form $\mathcal{Z}^{\mathsf{train}}$ are drawn at random, then how many groups do we need such that the affine hull captures ${\mathcal{Z}^{\times}}$, i.e., $\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})={\mathcal{Z}^{\times}}$, at which point CRM can achieve Cartesian Product Extrapolation (CPE). Another way to think about this is to say, how fast does the affine hull grow and capture the Cartesian product set ${\mathcal{Z}^{\times}}$? Next, we answer this question.
@@
Consider the the general setting with $m$ attributes, where each attribute takes $d$ possible values, leading to $d^{m}$ possible attribute combinations. Suppose we sample $s$ attribute vectors $z$ that comprise the support $\mathcal{Z}^{\mathsf{train}}$ uniformly at random (with replacement) from these $d^{m}$ possibilites. In the next theorem, we show that if the number of sampled attribute vectors exceeds $2c(md+d\log(d))$, then $\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})$ contains all the possible $d^{m}$ combinations with a high probability (greater than $1-\frac{1}{c}$), hence CRM achieves CPE. We want to emphasize this surprising finding: with almost a linear growth in $m$ and $d$, CRM generalizes to exponentially many $d^{m}$ groups.
@@
Consider the setting where $p(.|z)$ follows AED $\forall z\in{\mathcal{Z}^{\times}}$, $\mathcal{Z}^{\mathsf{train}}$ comprises of $s$ attribute vectors $z$ drawn uniformly at random from ${\mathcal{Z}^{\times}}$, and the test distribution $q$ satisfies compositional shift characterization with $\mathcal{Z}^{\mathsf{test}}=\mathcal{Z}^{\times}$. If $s\geq 2c(md+d\log(d))$, where $d$ is sufficiently large, $\hat{p}(z|x)=p(z|x),\forall z\in\mathcal{Z}^{\mathsf{train}},\forall x\in\mathbb{R}^{n}$, $\hat{q}(z)=q(z),\forall z\in{\mathcal{Z}^{\times}}$, then the output of CRM (equation 10) matches the test distribution, i.e., $\hat{q}(z|x)=q(z|x)$, $\forall z\in{\mathcal{Z}^{\times}},\forall x\in\mathbb{R}^{n}$, with probability greater than $1-\frac{1}{c}$.
In Section 10.3, we first present the proof for the case with $m=2$ attributes and provide visual illustrations to assist the reader. It is then followed by the more involved proof for the general case of $m$ attributes in Section 10.4.
3 Algorithm for CRM
Input: Training set $\mathcal{D}^{\mathrm{train}}=\{(x,z)\}$
Output: Classifier parameters $\hat{\theta}$, $\hat{W}$, $B^{\star}$
@@ 3 Algorithm for CRM
  • Estimate train prior $\hat{p}(z)$ based on group counts in $\mathcal{D}^{\mathrm{train}}$
  • Compute the energy terms as $\tilde{E}(x)=\tilde{W}\phi(x;\tilde{\theta})$
  • CRM Step 1: Train additive energy classifier $\tilde{p}$ (e.q. 7) by emprirical risk minimization:
    $\hat{\theta},\hat{W},\hat{B}\in\operatorname*{arg\,min}_{\tilde{\theta},\tilde{W},\tilde{B}}R(\tilde{p})$
  • CRM Step 2: Estimate extrapolated bias $B^{\star}$ (e.q. 11) via an average over training examples.
@@
  • Set $\hat{q}(z)$ as uniform prior over all groups.
  • Compute test group probabilities $\hat{q}(z|x)$ via e.q. 10 using $\hat{E}$ and $B^{\star}$ learned during training.
@@
In a nutshell, CRM consists of: a) training additive energy classifier $\hat{p}(z|x)$ (e.q. 7) by maximum likelihood (e.q.9) for trainset group prediction; b) compute extrapolated biases $B^{\star}$ (e.q.11); c) infer group probabilities on compositionally shifted test distribution using $\hat{q}(z|x)$ (e.q 10). Algorithm 1 provides the associated pseudo-code, where we have a basic architecture using a deep network backbone $\phi(x;\theta)$ followed by a linear mapping (matrix $W$)  [^1] Instead of linear, we could use separate non-linear heads to obtain the energy components for each attribute.. For the case where we have 2 attributes $z=(y,a)$ (illustrated in Figure 1), we provide a detailed algorithm and PyTorch implementation in Section 9.
@@ 4.1 Setup
We evaluate CRM on widely recognized benchmarks for subpopulation shifts (Yang et al., 2023b), that have attributes $z=(y,a)$, where $y$ denotes the class label and $a$ denotes the spurious attribute ($y$ and $a$ are correlated). However, the standard split between train and test data mandated in these benchmarks does not actually evaluate compositional generalization capabilities, because both train and test datasets contain all the groups ($\mathcal{Z}^{\mathsf{train}}=\mathcal{Z}^{\mathsf{test}}={\mathcal{Z}^{\times}}$). Therefore, we repurpose these benchmarks for compositional shifts by discarding samples from one of the groups ($z$) in the train (and validation) dataset; but we don’t change the test dataset, i.e., $z\not\in\mathcal{Z}^{\mathsf{train}}$ but $z\in\mathcal{Z}^{\mathsf{test}}$. Let us denote the data splits from the standard benchmarks as $({\mathcal{D}}_{\mathsf{train}},{\mathcal{D}}_{\mathsf{val}},{\mathcal{D}}_{\mathsf{test}})$. Then we generate multiple variants of compositional shifts $\{({\mathcal{D}}_{\mathsf{train}}^{\neg z},{\mathcal{D}}_{\mathsf{val}}^{\neg z},{\mathcal{D}}_{\mathsf{test}})\;|\;z\in{\mathcal{Z}^{\times}}\}$, where ${\mathcal{D}}_{\mathsf{train}}^{\neg z}$ and ${\mathcal{D}}_{\mathsf{val}}^{\neg z}$ are generated by discarding samples from ${\mathcal{D}}_{\mathsf{train}}$ and ${\mathcal{D}}_{\mathsf{val}}$ that belong to the group $z$.
@@
Following this procedure, we adapted Waterbirds (Wah et al., 2011), CelebA (Liu et al., 2015), MetaShift (Liang and Zou, 2022), MultiNLI (Williams et al., 2017), and CivilComments Borkan et al. (2019) for experiments. We also experiment with the NICO++ dataset (Zhang et al., 2023), where we already have $\mathcal{Z}^{\mathsf{train}}\subsetneq\mathcal{Z}^{\mathsf{test}}={\mathcal{Z}^{\times}}$ as some groups were not present in the train dataset. However, these groups are still present in the validation dataset ($\mathcal{Z}^{\mathsf{val}}={\mathcal{Z}^{\times}}$). Hence, the only transformation we apply to NICO++ is to drop samples from the validation dataset so that $\mathcal{Z}^{\mathsf{train}}=\mathcal{Z}^{\mathsf{val}}$. Note that our benchmarks cover diverse scenarios, with binary (Waterbirds, CelebA, MetaShift) and non-binary attributes (MultiNLI, CivilComments, NICO++), resulting in total groups varying from $4$ to $360$ (NICO++).
@@
For baselines, we train classifiers via ERM, Group Distributionally Robust Optimization (GroupDRO) (Sagawa et al., 2019), Logit Correction (LC) (Liu et al., 2022b), supervised logit adjustment (sLA) (Tsirigotis et al., 2024), Invariant Risk Minimization (IRM) (Arjovsky et al., 2019), Risk Extrapolation (VREx) (Krueger et al., 2021), and Mixup (Zhang et al., 2017). In all cases we employ a pretrained architecture as the representation network $\phi$, followed by a linear layer $W$ to get class predictions, and fine-tune them jointly (see Section 12.3 for details).
For evaluation metrics, we computed the average accuracy, group-balanced accuracy, and worst-group accuracy (WGA) on the test set. Due to imbalances in group distribution, a method can obtain good average accuracy despite having bad worst-group accuracy. Therefore, WGA is more indicative of robustness to spurious correlations (Section 12.2).
@@ 4.2 Results
Table 1 shows the results of our experiment, where due to space constraints, we only compare with the best performing baselines and don’t report the group balanced accuracy. Complete results with comparisons with IRM, VREx, and Mixup are provided in Appendix 13.1 (Table 5). For each dataset and metric, we report the average performance over its various compositional shift scenarios $\{({\mathcal{D}}_{\mathsf{train}}^{\neg z},{\mathcal{D}}_{\mathsf{val}}^{\neg z},{\mathcal{D}}_{\mathsf{test}})\;|\;z\in{\mathcal{Z}^{\times}}\}$ (detailed results for all scenarios are in Section 13.2). In all cases, CRM either outperforms or is competitive with the baselines in terms of worst group accuracy (WGA).
Further, for Waterbirds and MultiNLI, while the logit adjustment baselines appear competitive with CRM on average, if we look more closely at the worst case compositional shift scenario, we find these baselines fare much worse than CRM. For Waterbirds, LC obtains $69.0\%$ WGA while CRM obtains $73.0\%$ WGA for the worst case scenario of dropping the group $(0,1)$ (Table 6). Similarly, for MultiNLI, sLA obtains $19.7\%$ WGA while CRM obtains $31.0\%$ WGA for the worst case scenario of dropping the group $(0,0)$ (Table 9).
@@
We also report the worst group accuracy (other metrics in Table 14Section 13.5) for the original benchmark $({\mathcal{D}}_{\mathsf{train}},{\mathcal{D}}_{\mathsf{val}},{\mathcal{D}}_{\mathsf{train}})$, which was not transformed for compositional shifts, denoted WGA (No Groups Dropped). This can be interpreted as the “oracle” performance for that benchmark, and we can compare methods based on the performance drop in WGA due to discarding groups in compositional shifts. ERM and GroupDRO appear the most sensitive to compositional shifts, and the logit adjustment baselines also show a sharp drop for the CelebA benchmark; while CRM is more robust to compositional shifts.
Multiple spurious attributes case. We benchmark CRM for class label prediction with multiple spurious attributes. For this, we augment the CelebA benchmark to have three additional attributes (total $m=5$ attributes). We provide setup details in Section 13.3, and Table 12 presents the associated results. We find that CRM still continues to be the superior method (as per WGA), in fact, the difference in WGA between CRM and runner up baseline increases to $28.9\%$, as compared to $14\%$ for the case of CelebA 2-attribute (Table 2). Hence, the benefits with CRM become more pronounced for multiple spurious attributes settings.
@@
Importance of extrapolating the bias. We conduct an ablation study for CRM where we test a variant that uses the learned bias $\hat{B}$ (e.q. 9) instead of the extrapolated bias $B^{\star}$ (e.q. 11). Results are presented in Table 2. They show a significant drop in worst-group accuracy if we use the learned bias instead of the extrapolated one. Hence, our theoretically grounded bias extrapolation step is crucial to generalization under compositional shifts. In Section 13.4 (Table 13) we conduct further ablation studies, showing the impact of different choices of the test log prior.
Further, we analyze CRM’s performance with varying number of train groups, details in Section 13.6.
@@
Table 1: Robustness under compositional shift. We compare the proposed CRM method to baseline ERM classifier training with no group information, and to robust methods that leverage group labels: G-DRO, LC, and sLA. We report test Average Accuracy and Worst Group Accuracy (WGA) (mean over 3 random seeds), averaged as a group is dropped from training and validation sets. Last column is WGA under the dataset’s standard subpopulation shift benchmark, i.e. with no group dropped. All methods have a harder time to generalize when groups are absent from training, but CRM appears consistently more robust. Full results with balanced group accuracy, and baselines IRM, VREx, and Mixup in Appendix (Table 5).
@@
5 Related works
We briefly describe the relevant prior works, with a detailed discussion in Section 7.
Compositional Generalization Compositionality has long been seen as an essential capability (Fodor and Pylyshyn, 1988; Hinton, 1990; Plate et al., 1991; Montague, 1970) on the path to building human-level intelligence. The history of compositionality being too long to cover in detail here, we refer the reader to these surveys (Lin et al., 2023; Sinha et al., 2024). Most prior works have focused on generative aspect of compositionality, where the model needs to recombine individual distinct factors/concepts and generate the final output in the form of text (Gordon et al., 2019; Lake and Baroni, 2023) or image (Liu et al., 2022a; Wang et al., 2024). For image generation in particular, a fruitful line of work is rooted in additive energy based models (Du et al., 2020, 2021; Liu et al., 2021; Nie et al., 2021), which translates naturally to additive diffusion models (Liu et al., 2022a; Su et al., 2024). Our present work also leverages an additive energy form, but our focus is on learning classifiers robust under compositional shifts, rather than generative models.
Domain Generalization Generalization under subpopulation shifts, where certain groups or combinations of attributes are underrepresented in the training data, is a well-known challenge in machine learning. GroupDRO (Sagawa et al., 2019) is a prominent method that minimizes the worst-case group loss to improve robustness across groups. IRM (Arjovsky et al., 2019) encourages the model to learn invariant representations that perform well across multiple environments. Closely related to our proposed method are the logit adjustment methods, LC (Liu et al., 2022b) and sLA (Tsirigotis et al., 2024) that use logit adjustment for group robustness. There are many other interesting approaches that were proposed, see the survey (Zhou et al., 2022) for details. The theoretical guarantees developed for these approaches (Arjovsky et al., 2019; Rosenfeld et al., 2020; Ahuja et al., 2020) require a large diversity in terms of the environments seen at the training time. In our setting, we incorporate inductive biases based on additive energy distributions that help us arrive at provable generalization with limited diversity in the environments.
6 Conclusion
We provide a novel approach (CRM) based on flexible additive energy models for compositionality in discriminative tasks. CRM can provably extrapolate to novel attribute combinations within the discrete affine hull of the training support, where the affine hull grows quickly with the training groups to cover the Cartesian product extension of the training support. Our empirical results demonstrate that the additive energy assumption is sufficiently flexible to yield good classifiers for high-dimensional images, and CRM is able to extrapolate to novel combinations in $\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})$, without having to model high-dimensional $p(x|z)$ nor having to estimate their partition function. CRM is a simple and efficient algorithm that empirically proved consistently more robust to compositional shifts than approaches based on other logit-shifting schemes and GroupDRO.